Grade 12Physical Chemistry

What is the minimum volume of water required to dissolve 1g of calcium sulphate at 298 K? (For calcium sulphate, Ksp is 9.1 × 10–6). ?

sudhanshu
12 Years agoGrade 12

2 Answers

Gaurav
11 Years ago
Hello Student
Ksp = [Ca2+][SO42-]
Let the solubility is S
Then
Ksp = s2
9,1 x 10-6 = s2
S = 3.02 x 10-3 mol/L
Molecular mass of CuSO4 = 136 g
Solubility in Gram/L = 3.02 x 10-3 x 136 = 0.41
Therefore to dissolve 1 g of salt we require 1/0.41 = 2.44 L water
Yash Chourasiya
6 Years ago
Dear Student

Let s be the solubility ofCaSO4​.
CaSO4​ ⇌ Ca2+ SO42−
[Ca2+] = [SO42+​] =s
Ksp​[Ca2+][SO42−​] = s×s = s2 =9.1×10−6
s = 0.003016 M
The molar mass ofCaSO4​is40 + 32 + 64 = 136 g.
The solubility in g/L is0.003017 × 136 = 0.410 g/L.
Hence, 0.41 g dissolves in 1 L.
1 g will dissolve in 1/0.41 ​= 2.43 L

I hope this answer will help you.
Thanks & Regards
Yash Chourasiya